# fmeffects: An R Package for Forward Marginal Effects

### v1

@@ Abstract
Forward marginal effects (FMEs) have recently been introduced as a versatile and effective model-agnostic interpretation method. They provide comprehensible and actionable model explanations in the form of: If we change $x$ by an amount $h$, what is the change in predicted outcome $\widehat{y}$? We present the R package fmeffects, the first software implementation of FMEs. The relevant theoretical background, package functionality and handling, as well as the software design and options for future extensions are discussed in this paper.
@@ Introduction
Forward marginal effects (FMEs) (Scholbeck et al., 2022) provide simple yet accurate local model-agnostic explanations in terms of forward differences in prediction. They address questions of the form: If we change x by an amount $h$, what is the change in predicted outcome $\widehat{y}$? For instance, given a medical study where a model is trained to predict a patient’s disease risk, FMEs can tell us each patient’s individual change in predicted risk due to losing 5kg in body weight. FMEs thus provide actionable and comprehensible advice for stakeholders, including ones without expertise in machine learning. If the change in predicted risk is substantial enough, doctors may recommend a tailored exercise and nutrition regimen. To avoid interpreting the change in prediction (for instance, the change in predicted disease risk) as a linear effect, Scholbeck et al. (2022) developed a non-linearity measure (NLM) to supplement the FME. Furthermore, in applied statistics, FMEs are typically averaged to an average marginal effect (AME) for an entire data set, which serves as a global (scalar-valued) feature effect measure (Bartus, 2005). However, a single global average does not accurately represent the nuances of a non-linear model or one with complex interactions. As a remedy, Scholbeck et al. (2022) developed the conditional average marginal effect (cAME) estimator for subgroups of observations. They suggest to find subgroups with similar FMEs, for instance, by recursive partitioning (RP). FMEs therefore represent a means to explain models on a local, regional, and global level.
@@
Contributions: We present the R package fmeffects, the first software implementation of FMEs, including the NLM and the cAME. The user interface only requires a pre-trained model and an evaluation data set. The package is designed according to modular principles and thus, various parts of the package are easily maintainable and extensible with additional functionality. This paper introduces the relevant theoretical background, explains the handling of the package with an applied interpretation use case, and discusses the software design and options for future maintenance and extensions.
Related work: Marginal effects (MEs) have a long history in applied statistics and the Stata programming language (StataCorp, 2023). Initially implemented by Bartus (2005), the margins() command is now fully integrated into Stata and provides comprehensive capabilities for various computations and visualizations (Williams, 2012). MEs are typically defined in terms of derivatives of the model w.r.t. a feature. For instance, this variant is the default approach to interpret models in econometrics (Greene, 2012). A less commonly used definition uses forward differences in prediction with varying step sizes (Scholbeck et al., 2022; Mize et al., 2019). Note that - as opposed to forward differences - derivatives are not suitable to explain piecewise constant prediction functions such as tree-based models. In recent years, MEs have gained traction in the R community. The R package margins (Leeper, 2018) was the first port of Stata’s margins() command to R. Other packages related to MEs include ggeffects (Lüdecke, 2018) and marginaleffects (Arel-Bundock, 2023). Our package, fmeffects, mainly differs from such packages in two aspects:
  • Model-agnostic: Other R packages concerned with MEs directly implement support for specific models in R, which typically does not include model classes such as random forests, support vector machines, or neural networks. In contrast, we choose a model-agnostic approach. For one, we provide support for a plethora of models trained via mlr3 (Lang et al., 2019), tidymodels (Kuhn and Wickham, 2020), and caret (Kuhn, 2008); and second, we implement an adapter pattern (Gamma et al., 1994) (see section on Design and options for extensions) that allows extensions with new model classes without substantially interfering with the remaining source code.
  • Implementing new theory surrounding FMEs: The fmeffects package is the first complete software implementation of the theory surrounding novel model-agnostic FMEs as introduced by Scholbeck et al. (2022).
Methodological background
@@
We first provide a brief overview of the relevant theory. For an in-depth introduction, we refer the reader to the theoretical paper by Scholbeck et al. (2022). FMEs can be used for model explanations on the local, regional (also referred to as semi-global), and global level. These differ with respect to the scope of the feature space where the explanation applies. The local level explains models for single data points; the regional level for a subspace of the entire feature space (which, for randomly sampled data, translates to subgroups of observations), and the global level for the entire feature space (which, for randomly sampled data, translates to the entire data set). Increasing the “topographical” scope of the explanation requires increasing amounts of aggregations of local explanations (see for instance, the illustration by Scholbeck et al. (2020) of aggregations of local explanations to global ones for various methods). This can be problematic for non-linear models where local explanations can be highly heterogeneous due to interactions and non-linear effects estimated by the model.
@@
Forward marginal effects
The FME can be considered a basic, local unit of interpretation. Given an observation $\bm{x}$, it tells us how the prediction changes if we move $\bm{x}$ to a different location in feature space $\bm{x}^{\ast}$. This change in feature values typically only concerns a subset of features S, where $\bm{x}^{\ast}=(\bm{x}_{S}+\bm{h}_{S},\bm{x_{-S}})$:
For categorical features, this shift is only possible for observations whose observed values differ from the reference location. The observation-specific categorical FME with reference category $c_{j}$ corresponds to:
Non-linearity measure
For continuous features, we can consider $\bm{x}_{S}+\bm{h}_{S}$ a continuous transition of feature values. The associated change in prediction may be misinterpreted as a piecewise linear effect, which is counteracted by the non-linearity measure (NLM). The NLM corresponds to a continuous coefficient of determination $R^{2}$ between prediction function and the linear secant that intersects $\bm{x}$ and $\bm{x}^{\ast}$ (see Fig. 1). We first parameterize this continuous transition through the feature space as a fraction $t\in[0,1]$ of the multivariate step size $\bm{h}_{S}$:
$\gamma(t)=\bm{x}+t\cdot(\bm{h}_{S},0,\dots,0)$
Given the linear secant $g_{\bm{x},\bm{h}_{S}}(\gamma(t))$, the NLM corresponds to:
@@
Fig. 2 illustrates the setting for multivariate feature changes. The NLM can be approximated via numerical integration, for instance via Simpson’s rule.
Figure 1: Illustration by Scholbeck et al. (2022) of a univariate FME (blue bar) given the prediction function (yellow) and linear secant (green). The NLM indicates how well the secant can explain the prediction function - which is inversely proportional to the integral visualized in red - in relation to how well the most uninformative baseline model - the average prediction - can explain the prediction function.
Figure 2: Illustration of the multivariate NLM by Scholbeck et al. (2022). Left: An exemplary bivariate prediction function and two points to compute an FME. Our observation is located at point (-5, -5), and we create the shortest path through the feature space to reach point (5, 5), which consists of directly proportional changes in both features. Above the path, we see the linear secant and the non-linear prediction function. Right: The multivariate change in feature values can be parameterized as a percentage $t$ of the step size $\bm{h}_{S}$. As a result, the deviation between prediction function and linear secant, as well as the deviation between the prediction function and mean prediction both correspond to a line integral which can be numerically approximated, for instance via Simpson’s rule.
The NLM indicates how well the linear secant is able to explain the prediction function, compared to the baseline model of using the mean prediction. A value of 1 indicates perfect linearity, where the linear secant is identical to the prediction function. For a value of 0, the mean prediction is able to explain the prediction function as equally well as the secant. For negative values, the mean prediction better explains the prediction function than the linear secant (severe non-linearity).
@@
In order to receive a global model explanation akin to a beta coefficient in linear models, local FMEs can be averaged to a global average marginal effect (AME). However, this does not accurately represent the nuances of a non-linear prediction function. Given a data set $\mathcal{D}$, the conditional average marginal effect (cAME) estimator applies to a subgroup of observations $\mathcal{D}_{[\;]}$:
@@
$=\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{f}\left(\bm{x_{S}}^{(i)}+\bm{h}_{S},\bm{x_{-S}}^{(i)}\right)-\widehat{f}\left(\bm{x^{(i)}}\right)\right)\,\forall i:\bm{x^{(i)}}\in\mathcal{D}_{[\;]}$
@@
Although this estimator can be applied to arbitrary subgroups, an interpretation task demands searching for subgroups with homogeneous local model explanations. Desiderata for such subgroups include within-group effect homogeneity, between-group effect heterogeneity, full segmentation, non-congruence, confidence, and stability. A viable option to partition $\mathcal{D}$ is to run a recursive partitioning algorithm on $\mathcal{D}$ with FMEs as the target which aims at finding subgroups where FMEs are similar. For instance, in fmeffects, both rpart (Therneau and Atkinson, 2019) and ctree() from partykit (Hothorn and Zeileis, 2015) are supported to find subgroups.
@@
The fme() function is the central user interface, which - apart from certain parameters - only requires a pre-trained model and a data set (see section Design and options for extensions for details). Further arguments specify the target, feature(s), step size(s), whether to compute NLM values for each FME, and specification of an extrapolation detection method [^3] Model extrapolations are predictions in areas with a low density of training observations. Scholbeck et al. (2022) provide a short introduction to this topic.. The fme() function initiates construction of and computations for a ForwardMarginalEffects object without requiring the user to know R6 (Chang, 2021) syntax.
For this use case, we train a random forest from the ranger package (Wright and Ziegler, 2017) on the bike sharing data set (Fanaee-T, 2013) using mlr3. Note that models trained via tidymodels and caret are also supported. Our goal is to predict and explain the daily bike rental demand in Washington D.C. based on features such as the outside temperature, wind speed, or humidity. We first train the model:
@@
Then, we simply pass the trained model, evaluation data, and remaining parameters to the fme() function. It returns a ForwardMarginalEffects object which can be analyzed via summary() and visualized via plot() (see Fig. 3). Here, the outside temperature is raised by 5 degrees Celsius ceteris paribus. Each dot represents a single instance. The arrow above the horizontal axis indicates the step size and direction of the feature change in relation to the feature range. The horizontal value of each instance indicates its observed feature value in the provided data set; the vertical axis value is the corresponding FME.
@@
Let us single out the observation with the largest associated FME (brown colored and enlarged in Fig. 3). This observation corresponds to a single day with a recorded temperature of 13.12 degrees Celsius. Increasing the temperature by 5 degrees Celsius on this particular day results in an additional 125 predicted bike rentals. We plot such model explanations for the entire data set and average FMEs to receive a global model explanation. The AME - the global average of FMEs - is 14.5: An increase in temperature by 5 degrees Celsius results in an average change in predicted daily bike rentals by 14.5.
@@
> effects = fme(+   model = forest,+   data = bikes,+   target = "count",+   feature = "temp",+   step.size = 5,+   ep.method = "envelope")> summary(effects)
@@
Forward Marginal Effects ObjectStep type:  numericalFeatures & step lengths:  temp, 5Extrapolation point detection:  envelope, EPs: 81 of 727 obs. (11 %)Average Marginal Effect (AME):  14.4844
> plot(effects)
@@
Figure 3: Plot of univariate FMEs for feature ‘temp’ and step size 5. Each dot represents an observation. The horizontal value represents the observed value of feature ‘temp’. We move each observation’s ‘temp’ value according to the arrow’s direction and length (in this case, 5 units to the right). The respective FME value is indicated by the vertical value of each point. The horizontal bar indicates the AME (the global FME average). An additional smoothing function facilitates interpretations by modeling an approximate pattern of FMEs across the feature range. We single out one observation (brown colored and enlarged) to demonstrate a local interpretation.

### v2

@@ Abstract
Forward marginal effects have recently been introduced as a versatile and effective model-agnostic interpretation method particularly suited for non-linear and non-parametric prediction models. They provide comprehensible model explanations of the form: if we change feature values by a pre-specified step size, what is the change in the predicted outcome? We present the R package fmeffects, the first software implementation of the theory surrounding forward marginal effects. The relevant theoretical background, package functionality and handling, as well as the software design and options for future extensions are discussed in this paper.
@@ Introduction
A growing number of disciplines are adopting black box machine learning (ML) models to make predictions, including medicine (Rajkomar et al., 2019; Boulesteix et al., 2020), psychology (Dwyer et al., 2018), economics (Mullainathan and Spiess, 2017; Athey and Imbens, 2019), or the earth sciences (Dueben and Bauer, 2018). Although one can often observe a superior predictive performance of black box models (such as neural networks, gradient boosting, random forests, or support vector machines) over intrinsically interpretable models (such as generalized linear or additive models), their lack of transparency or interpretability is considered a major drawback (Breiman, 2001). This has been a major driver in the development of model-agnostic explanation techniques, which are often referred to by the umbrella terms of interpretable ML (Molnar, 2022) or explainable artificial intelligence (Kamath and Liu, 2021).
Marginal effects (MEs) (Williams, 2012) have been a mainstay of model interpretations in many applied fields such as econometrics (Greene, 2019), psychology (McCabe et al., 2022), or medical research (Onukwugha et al., 2015). MEs explain the effect of features on the predicted outcome in terms of derivatives w.r.t. a feature or forward differences in prediction. They are typically averaged to an average marginal effect (AME) for an entire data set, which serves as a global (scalar-valued) feature effect measure (Bartus, 2005). To explain feature effects for non-linear models, Scholbeck et al. (2024) introduced a unified definition of forward marginal effects (FMEs), a non-linearity measure (NLM) for FMEs, and the conditional average marginal effect (cAME). The NLM is an auxiliary model diagnostic to avoid interpreting local changes in prediction as linear effects. The cAME aims to describe the model via regional FME averages for subgroups with similar FMEs, which can, for instance, be found by recursive partitioning (RP). FMEs, therefore, represent a means to explain models on a local, regional, and global level.
@@
Contributions: We present the R package fmeffects, the first software implementation of the theory surrounding FMEs, including the NLM and the cAME. The user interface only requires a pre-trained model and an evaluation data set. The package is designed according to modular principles, making it simple to maintain and extend. This paper introduces the relevant theoretical background of FMEs, demonstrates the usage of the package in the context of a practical use case, and explains the software design.
Background on forward marginal effects
@@
FMEs can be used for model explanations on the local, regional (also referred to as semi-global), and global level. These differ with respect to the region of the feature space that the explanation refers to. The local level explains a model/prediction for single observations, the regional level for a certain subspace (or subgroups of observations), and the global level for the entire feature space. Increasing the scope of the explanation requires increasing amounts of aggregations of local explanations (see the illustration by Scholbeck et al. (2020) of aggregations of local explanations to global ones for various methods). This can be problematic for non-parametric models where local explanations can be highly heterogeneous due to non-linear effects or interactions.
@@
Let $\widehat{f}:\mathcal{X}\rightarrow\mathbb{R}$ be the prediction function of a learned model where $\mathcal{X}\subset\mathbb{R}^{p}$ denotes the feature space. While our definition naturally covers regression models, for classification models, we assume that $\widehat{f}$ returns the score or probability for a predefined class of interest. A subspace of the feature space is denoted by $\mathcal{X}_{[\;]}\subseteq\mathcal{X}$. The random feature vector is denoted by [^1] Bold letters denote vectors. $\bm{X}=(X_{1},\dots,X_{p})$. Observations are denoted by $\bm{x}=(x_{1},\dots,x_{p})\in\mathcal{X}$. A set of feature indices is denoted by $S\subseteq\{1,\dots,p\}$. We often index (random) vectors as $\bm{x}_{S}$ or $\bm{X}_{S}$. We denote set complements by $-S=\{1,\,\dots\,,\,p\}\;\setminus\;S$. With slight abuse of notation, we represent the partitioning of a vector into two arbitrary but disjoint groups by $\bm{x}=(\bm{x}_{S},\bm{x}_{-S})$, regardless of the order of features. For a single feature of interest, the set $S$ is replaced by an integer index $j$. We usually assume an evaluation data set $\mathcal{D}=\left(\bm{x}^{(i)}\right)_{i=1}^{n}$, with $\bm{x}^{(i)}\in\mathcal{X}$, which may consist of both training and test data.
Forward marginal effects
The FME can be considered a basic, local unit of interpretation. Given an observation $\bm{x}$, it tells us how the prediction changes if we change a subset of feature values $\bm{x}_{S}$ by a vector of step sizes $\bm{h}_{S}$.
Scholbeck et al. (2024) introduced an observation-specific categorical FME whose definition is congruent with the FME for continuous features. The categorical FME corresponds to the change in prediction when replacing $x_{j}$ by the reference category $c_{j}$:
Note that this definition of a categorical ME differs from the one that is typically found in fields like econometrics (Williams, 2012), where we set $x_{j}$ to a reference category for all observations and then record the change in prediction resulting from changing the reference category to another category.
Furthermore, it is common to globally average MEs to an average marginal effect (AME) to estimate the expected local effect. For FMEs, this corresponds to:
$\;\text{AME}_{\mathcal{D},\bm{h}_{S}}$ $=\widehat{\mathbb{E}_{\bm{X}}\left[\text{FME}_{\bm{X},\bm{h}_{S}}\right]}$
$=\frac{1}{n}\sum_{i=1}^{n}\left[\widehat{f}\left(\bm{x}_{S}^{(i)}+\bm{h}_{S},\bm{x}_{-S}^{(i)}\right)-\widehat{f}\left(\bm{x}^{(i)}\right)\right]$ (1)
Note that for categorical feature changes and observations where $x_{j}=c_{j}$, the FME equals 0. In the fmeffects package, the categorical AME only consists of observations whose observed feature value differs from the selected category. This approach is motivated by our goal to explain the effects of changing feature values on the predicted outcome. For instance, in Fig. 11, we demonstrate the effect of rainfall on the daily number of bike rentals in Washington D.C. by switching each non-rainy day’s precipitation status to rainfall. Considering all observations, including rainy days, would obfuscate the interpretation we desire from our model. However, it is important to remember that every AME comprises a different set of points.
Step size selection
The selection of step sizes is determined by contextual and data-related considerations (Scholbeck et al., 2024). First, the FME allows us to investigate the model according to specific research questions. For instance, we might be interested in the effects of a specific change in a patient’s body weight on the predicted individual disease risk. Often, we are interested in an interpretable or intuitive step size. In the case of body weight, typically expressed in kilograms, we could use a 1kg change (for instance, instead of 1g) as a natural increment. Without contextual information, we could use a unit change as a reasonable default; or dispersion-based measures such as one standard deviation, percentages of the interquartile range, or the mean/median absolute deviation.
Non-linearity measure
For continuous features, we can consider $\bm{x}_{S}+\bm{h}_{S}$ a continuous transition of feature values. The associated change in prediction may be misinterpreted as a linear effect. This is counteracted by the NLM, which corresponds to a continuous coefficient of determination $R^{2}$ between the prediction function and the linear secant that intersects $\bm{x}$ and $(\bm{x}_{S}+\bm{h}_{S},\bm{x}_{-S})$ (see Fig. 1). The continuous transition through the feature space is first parameterized as a fraction $t\in[0,1]$ of the multivariate step size $\bm{h}_{S}$:
$\gamma(t)=\begin{pmatrix}x_{1}\\ \vdots\\ x_{p}\end{pmatrix}+t\cdot\begin{pmatrix}h_{1}\\ \vdots\\ h_{s}\\ 0\\ \vdots\\ 0\end{pmatrix}\quad,\quad t\in[0,1]$
The value of the linear secant $g_{\bm{x},\bm{h}_{S}}(t)$ corresponds to:
@@
The mean prediction $\widehat{f}_{\text{mean}}$ on the interval $t\in[0,1]$ is given by:
@@
[…]
The NLM compares the squared deviation between the prediction function and the linear secant to the squared deviation between the prediction function and the mean prediction:
[…]
Fig. 2 illustrates the setting for multivariate feature changes. The NLM can be approximated via numerical integration, e.g., via Simpson’s rule.
Figure 1: Illustration by Scholbeck et al. (2024) of a univariate FME (blue) given the prediction function (black) and linear secant (orange, dashed). The NLM indicates how well the secant can explain the prediction function (inversely proportional to the purple area) compared to how well the most uninformative baseline model (the average prediction) can explain the prediction function.
Figure 2: Illustration of the multivariate NLM by Scholbeck et al. (2024). Left: An exemplary bivariate prediction function and two points to compute an FME. Consider an observation $\bm{x}=(-5,-5)$ and step size vector $\bm{h}_{S}=(10,10)$. We create the shortest path through the feature space to reach the point (5, 5), which consists of directly proportional changes in both features. Above the path, we see the linear secant (orange, dashed) and the non-linear prediction function (black). Right: The multivariate change in feature values can be parameterized as a percentage $t$ of the step size $\bm{h}_{S}$. The deviation between the prediction function and the linear secant, as well as the deviation between the prediction function and mean prediction, both correspond to a line integral.
The NLM indicates how well the linear secant can explain the prediction function, compared to the baseline model of using the mean prediction. A value of 1 indicates perfect linearity, where the linear secant is identical to the prediction function. For a value of 0, the mean prediction can explain the prediction function to the same degree as the secant. For negative values, the mean prediction better explains the prediction function than the linear secant (severe non-linearity).
@@
To receive a global model explanation akin to a beta coefficient in linear models, local FMEs can be averaged to the AME. Mehrabi et al. (2021) define an aggregation bias as drawing false conclusions about individuals from observing the entire population. Given a data set $\mathcal{D}$, the conditional average marginal effect (cAME) estimator applies to a subgroup of $n_{[\;]}$ observations, denoted by $\mathcal{D}_{[\;]}$:
@@
$=\frac{1}{n_{[\;]}}\sum_{i\,:\,\bm{x}^{(i)}\in\mathcal{D}_{[\;]}}\left[\widehat{f}\left(\bm{x}_{S}^{(i)}+\bm{h}_{S},\bm{x}_{-S}^{(i)}\right)-\widehat{f}\left(\bm{x}^{(i)}\right)\right]$ (2)
@@
Although this estimator can be applied to arbitrary subgroups, we aim to find subgroups with cAMEs that counteract the aggregation bias. Desiderata for such subgroups include within-group effect homogeneity, between-group effect heterogeneity, full segmentation, non-congruence, confidence, and stability (Scholbeck et al., 2024). In other words, we aim to partition the data into subgroups that explain variability in the FMEs. A viable option to partition $\mathcal{D}$ is to run RP on $\mathcal{D}$ with FMEs as the target. For instance, in fmeffects, both rpart (Therneau and Atkinson, 2019) and ctree() from partykit (Hothorn and Zeileis, 2015) are supported to find subgroups.
Related work
Model-agnostic interpretations
The basic mechanism behind model-agnostic methods is to probe the model with different feature values, a methodology similar to a model sensitivity analysis (Scholbeck et al., 2020, 2023). The basis of explaining models is to determine the direction and magnitude of the effect of features on the predicted outcome (Casalicchio et al., 2019; Scholbeck et al., 2020, 2024). The individual conditional expectation (ICE) (Goldstein et al., 2015), partial dependence (PD) (Friedman, 2001), accumulated local effects (ALE) (Apley and Zhu, 2020), Shapley values (Štrumbelj and Kononenko, 2010; Lundberg and Lee, 2017; Covert et al., 2020) and local interpretable model-agnostic explanations (LIME) (Ribeiro et al., 2016) are some of the most popular model-agnostic explanation methods for ML models. Notably, counterfactual explanations (Wachter et al., 2018) represent the reverse of the FME, indicating the smallest necessary change in feature values to reach a targeted prediction.
FMEs complement the literature by allowing for a unique combination of local, regional, and global model explanations. Furthermore, while most methods (including the ICE, PD, ALE, or Shapley values) provide explanations in terms of prediction levels, FMEs provide explanations in terms of prediction changes. LIME is based on training a local and interpretable surrogate model whose coefficients can also provide an interpretation in terms of prediction changes. Scholbeck et al. (2024) highlighted differences between both approaches: notably, while surrogate models introduce additional uncertainty connected with the estimation of the surrogate, FMEs are motivated by the goal of stable and comprehensible model insight. Furthermore, locally estimated FMEs can be aggregated within subgroups and entire data sets for regional and global explanations. Around the same time, regional aggregations have also been introduced for ICE curves, for example (Britton, 2019; Herbinger et al., 2022; Molnar et al., 2024).
Relationship between individual conditional expectation and forward marginal effect
Scholbeck et al. (2024) illustrated a relationship between the ICE / PD and the FME / AME. In general, the FME can be seen as the difference between two locations on an ICE. The AME corresponds to the difference between two locations on the PD only for a function that is linear in the feature of interest. Therefore, the following relationship between the ICE and FME is worth noting here. The ICE can be considered a one-way sensitivity function that indicates the effects of varying a set of features indexed by $S$ while the remaining ones are kept constant:
$\text{ICE}_{\bm{x},S}(\bm{x}^{\ast}_{S})=\widehat{f}(\bm{x}^{\ast}_{S},\bm{x}_{-S})$
For an instance $\bm{x}$, the prediction after increasing $\bm{x}_{S}$ by $\bm{h}_{S}$ is also a value of the ICE:
$\text{FME}_{\bm{x},\bm{h}_{S}}$ $=\widehat{f}(\bm{x}_{S}+\bm{h}_{S},\bm{x}_{-S})-\widehat{f}(\bm{x})$
$=\text{ICE}_{\bm{x},S}(\bm{x}_{S}+\bm{h}_{S})-\text{ICE}_{\bm{x},S}(\bm{x}_{S})$
Related work on marginal effects
MEs have a long history in applied statistics and the Stata programming language (StataCorp, 2023). Initially implemented by Bartus (2005), the margins() command is now fully integrated into Stata and provides comprehensive capabilities for various computations and visualizations of statistical models such as (generalized) linear models (Williams, 2012). MEs are typically defined in terms of derivatives of the model w.r.t. a feature. For instance, this variant is the default approach to interpret models in econometrics (Greene, 2019). The FME is the less commonly used definition (Scholbeck et al., 2024; Mize et al., 2019). Note that—in contrast to forward differences—derivatives are not suitable to explain piecewise constant prediction functions such as tree-based models.
In recent years, MEs have gained traction in the R community. The R package margins (Leeper, 2018) was the first port of Stata’s margins() command to R. Other packages related to MEs include ggeffects (Lüdecke, 2018) and marginaleffects (Arel-Bundock, 2023). In particular, marginaleffects can also return FMEs (although under different terminology). Our package, fmeffects, mainly differs from marginaleffects in two aspects:
Implementing new theory surrounding FMEs:
The fmeffects package is the first software implementation of the theory surrounding model-agnostic FMEs as introduced by Scholbeck et al. (2024). Although packages such as marginaleffects support the computation of FMEs and other quantities, fmeffects is specifically designed for FMEs with unique features such as implementations of the NLM, the cAME via RP, and novel visualization methods.
Model-agnostic black box interpretations:
It follows that fmeffects is targeted at model-agnostic explanations of non-linear or intransparent models. Whereas existing theory on MEs (and packages such as marginaleffects) focuses on classical statistical modeling in combination with statistical inference (see, for instance, Breiman (2001) comparing statistical modeling culture with ML), FMEs (and thus fmeffects) are comparable to methods and software from the literature on interpretable ML such as the ICE, PD, ALE, or LIME. This does not imply that marginaleffects cannot be used for black box interpretations. As mentioned in the previous point, it also supports the computation of FMEs, e.g., in combination with mlr3, but the focus of fmeffects lies on the interpretation of black box models through a specialized and targeted range of novel capabilities.
Advantages and limitations of forward marginal effects
Advantages
Although the ICE and the FME are closely related, the latter provides several novel ways to generate insights into the model:
  • Univariate changes in feature values: FMEs are comparable to ICE curves for univariate changes in feature values. In certain scenarios, however, they may provide more comprehensible visualizations of effects for individual instances (see Fig. 4 for an example).
  • Bivariate changes in feature values: The ICE and PD also provide insight into the sensitivity of the model prediction for variations in two features, which is visualized as a heatmap (see Fig. 7). However, it is difficult to visually compare the ICE of many different observations (which correspond to heatmaps as well). Although the ICE provides insight into a larger variation in feature values, while the FME only considers a single tuple of changes in feature values, bivariate FMEs can be easily compared visually (see Fig. 6).
  • Higher-order changes in feature values: If we evaluate the sensitivity of the prediction for changes in more than two feature values, virtually every visualization method breaks down. In this case, FMEs still provide comprehensible model explanations that can be aggregated in various ways (see Fig. 10).
  • Local fidelity assessment: The locally restricted change in feature values for the FME facilitates evaluations of the fidelity of the model explanation (e.g., via the NLM). In other words, the NLM allows us to describe how well the FME summarizes the local shape of the prediction function in a single value. See Fig. 8 for a visualization of NLM values for different observations.
  • Comprehensible regional explanations: Although regional explanations have been first proposed in the context of grouping ICE curves (Herbinger et al., 2022; Britton, 2019), they more easily apply to scalar model explanations such as FMEs. Essentially, a regional model explanation represents a group of observations or a subspace of the feature space where model explanations are relatively homogeneous. Such groupings are easily achievable via RP or other techniques that do not require functional target values such as ICEs.
  • Avoiding extrapolation: The ICE is computed on the entire feature range (see, e.g., Fig. 4), which is likely to result in model extrapolations. By its nature, the FME is typically used with small step sizes relative to the feature range, which naturally avoids model extrapolations.
Limitations
  • Step size selection: The step size fundamentally influences effects and the model interpretation. Although FMEs for different step sizes can be computed and visualized in an exploratory manner, some level of prior reasoning about what step sizes to use is recommended.
  • Decision tree instability for cAME: Although not a shortcoming of the FME itself, subgroups found by RP to compute cAMEs are subject to a high variance. This may be counteracted by stabilizing the split search, e.g., by considering statistical significance of tree splits or resorting to different algorithms to find subgroups.
  • Non-linearity assessment for proportional feature changes: For multi-dimensional feature changes, the NLM only considers equally proportional changes in all features.
On causal interpretations and avoiding model extrapolations
Note that model-agnostic techniques, including FMEs, explain associations between the target and the features within the model. In the absence of additional assumptions, such associations cannot be interpreted as causes and effects (Molnar et al., 2022). For instance, increasing the value of a feature $x_{1}$ may always be accompanied by an increase in the target, but it may be the target $y$ that causes $x_{1}$ to increase. Another typical scenario is the presence of confounding factors that influence both $y$ and $x_{1}$. Finally, $x_{1}$ may only (or also) influence a mediator $x_{2}$, which in turn influences $y$.
This does not, however, make model interpretations obsolete. More importantly, as highlighted by Adadi and Berrada (2018), model interpretations can be used to gain knowledge, debug, audit, or justify the model and its predictions. Throughout this paper, we will model the effects of environmental influences on the number of daily bike rentals in Washington, D.C. For our estimated model, a drop in humidity by 10 percentage points has a considerable effect on the predicted number of daily bike rentals (see Fig. 5). This effect cannot be assumed to be causal, as humidity is physically influenced by the outside temperature, which will also affect people’s choice to rent a bike. Here, temperature is a confounder that influences both humidity and daily bike rentals. However, the business renting out bikes can still use the associations found by a model with a good predictive performance to control the optimal number of bikes at their disposal. This is conditional on the model’s ability to accurately predict the target for the given feature vector, requiring us to avoid model extrapolations, which correspond to predictions within areas of the feature space where the model has not seen much or any training data. This issue is closely linked to the multivariate distribution of the training data; in our example, a change in humidity is likely to be accompanied by a change in temperature as well, which we somewhat circumvent (depending on the magnitude of the step size) when making isolated changes to humidity. One may disregard this issue and deliberately predict in areas of the feature space the model has not seen during training. The resulting FMEs will still be valid model descriptions but, as explained above, they are likely to be bad descriptions of the data generating process.
Model extrapolations negatively impact many model-agnostic interpretation methods (Hooker, 2004b, a, 2007; Hooker et al., 2021; Molnar et al., 2022). For example, Apley and Zhu (2020) demonstrated how PD plots suffer from extrapolation issues and introduced ALE plots as a solution to this problem. Scholbeck et al. (2024) illustrated the perils of model extrapolations for FMEs specifically and discussed possible options. One option in particular is also implemented in fmeffects: points outside the multivariate envelope (meaning the Cartesian product of all observed feature ranges) of the training data can be excluded from the analysis. This directly relates to the selection of small step sizes relative to the feature range, as large step sizes will result in a point falling outside the envelope.
When using extrapolation prevention methods, note that we consider different sets of points for different step sizes, which differs from the usage of MEs in other contexts (see, for instance, the package marginaleffects for a comparison). The exclusion of points only impacts aggregations of FMEs, i.e., the cAME and AME. As discussed in the section on Forward marginal effects, this also affects the computation of categorical AMEs. In Eq. (1) and Eq. (2), the AME and cAME are formulated as estimators of the expected global or regional (concerning a subspace) effects. The fewer observations we are considering for an average, the larger the variance of the estimate.
@@
The fme() function is the central user interface. It mainly requires a pre-trained model and a data set (see section Design and options for extensions for details). Further control parameters include a list of features and step sizes, whether to compute NLM values for each FME, and an extrapolation detection method. The fme() function initiates the construction and computations of a ForwardMarginalEffect object without requiring the user to know R6 (Chang, 2021) syntax.
For this use case, we train a random forest from the ranger package (Wright and Ziegler, 2017) on the bike sharing data set (Fanaee-T, 2013) using mlr3. Note that models trained via tidymodels and caret are also supported, as well as models trained via lm(), glm(), and gam(). We aim to predict and explain the daily bike rental demand in Washington, D.C., based on features such as the outside temperature, wind speed, or humidity. We first train the model:
@@
Then, we simply pass the trained model, evaluation data, and remaining parameters to the fme() function. It returns a ForwardMarginalEffect object, which can be analyzed via summary() and visualized via plot() (see Fig. 3). Here, the outside temperature is raised by 5 degrees Celsius ceteris paribus. To avoid overplotting values, each hexagon represents a local average of FMEs. Users can easily access the data used by all plot functions to implement their own visualizations.
@@
Let us single out the observation with the largest associated FME. This observation corresponds to a single day with a recorded temperature of 8 degrees Celsius. Increasing the temperature by 5 degrees Celsius on this particular day results in 2699 additional predicted bike rentals. We plot such model explanations for the entire data set and average FMEs to receive a global model explanation. The AME—the global average of FMEs—is 307: an increase in temperature by 5 degrees Celsius results in an average increase of 307 predicted daily bike rentals.
@@
> effects.univariate.temp = fme(+   model = forest,+   data = bikes,+   features = list("temp" = 5),+   ep.method = "envelope")> summary(effects.univariate.temp)
@@
Forward Marginal Effects ObjectStep type:  numericalFeatures & step lengths:  temp, 5Extrapolation point detection:  envelope, EPs: 48 of 731 obs. (7 %)Average Marginal Effect (AME):  307.3275> plot(effects.univariate.temp)
@@
Figure 3: Plot of univariate FMEs for feature ‘temp’ and step size 5. Each hexagon represents a local FME average. The horizontal value represents the observed feature value of ‘temp’. Each observation’s ‘temp’ value is moved according to the arrow’s direction and length. The vertical value of each hexagon indicates the FME value associated with that feature change. The horizontal bar indicates the AME. The shade of the hexagon implies how many observations it contains. A smoothing function facilitates interpretations by modeling an approximate pattern of FMEs across the feature range.